Please answer parts d e and f. The answers to the rest have been answered repeatedly here and no one completes the problem. Thank you.To get information about the quality of the applicant pool, you pick a number 0 <r <n, and you callously) interview the first
Answer #1

d) Observe that

(pr Pr-1) n 1 1 r - i 1 i- 1 i=r+1 n 1 1 + i 1 - 1 n i-r+1 n 1 X n r1 i 1

Therefore,

1 1 1 2 7 2=r-1 X 2 1 1] 2 1 1 1 < 0 n(r 1 1 7 X X X

Hence qr is a strictly decreasing function of r for r > 0.

e) Observe that if pr is maximized at r, then,

i>0 Vi < >Pi-1 Vi <r

Hence, the result follows.

f) The following table gives us the value of qr at r:

r q
2 0.0828968254
3 0.0328968254
4 -0.0004365079365
5 -0.02543650794
6 -0.04543650794
7 -0.0621031746
8 -0.07638888889
9 -0.08888888889
10 -0.1

Hence the maximum for pr occurs at r = 3 and the value of the maximum is 0.398